### v1

# Adjacency spectral radius and $H$-factors in 1-binding graphs

Abstract
Let $G$ be a graph, and let $H:V(G)\longrightarrow\{\{1\},\{0,2\}\}$ be a set-valued function. Hence, $H(v)$ equals $\{1\}$ or $\{0,2\}$ for any $v\in V(G)$. We let
$H^{-1}(1)=\{v:v\in V(G)\ \mbox{and}\ H(v)=1\}.$
An $H$-factor of $G$ is a spanning subgraph $F$ of $G$ such that $d_{F}(v)\in H(v)$ for each $v\in V(G)$. Lu and Kano showed a characterization for the existence of an $H$-factor in a graph [Characterization of 1-tough graphs using factors, Discrete Math. 343 (2020) 111901]. Let $A(G)$ and $\rho(G)$ denote the adjacency matrix and the adjacency spectral radius of $G$, respectively. By using Lu and Kano’s result, we pose a sufficient condition with respect to the adjacency spectral radius to guarantee the existence of an $H$-factor in a 1-binding graph. In this paper, we prove that if a connected 1-binding graph $G$ of order $n\geq 11$ satisfies $\rho(G)\geq\rho(K_{1}\vee(K_{n-4}\cup K_{2}\cup K_{1}))$, then $G$ has an $H$-factor for each $H:V(G)\longrightarrow\{\{1\},\{0,2\}\}$ with $H^{-1}(1)$ even, unless $G=K_{1}\vee(K_{n-4}\cup K_{2}\cup K_{1})$.
Keywords: graph; binding number; adjacency spectral radius; $H$-factor.
(2020) Mathematics Subject Classification: 05C50, 05C70
1 Introduction
Throughout this paper, we only consider finite and undirected graphs without loops or multiple edges. Let $G$ denote a graph with vertex set $V(G)$ and edge set $E(G)$. For $v\in V(G)$, we let $d_{G}(v)$ and $N_{G}(v)$ denote the degree and the neighborhood of $v$ in $G$, respectively. Given $S\subseteq V(G)$, let $N_{G}(S)$ and $G[S]$ denote the set of neighbors of $S$ in $G$ and the subgraph of $G$ induced by $S$, respectively. Write $G-S=G[V(G)\setminus S]$. The number of isolated vertices and the number of components in $G$ are denoted by $i(G)$ and $\omega(G)$, respectively. The complete graph of order $n$ is denoted by $K_{n}$.
For vertex-disjoint graphs $G_{1}$ and $G_{2}$, the graph with vertex set $V(G_{1})\cup V(G_{2})$ and edge set $E(G_{1})\cup E(G_{2})$ is called the union of $G_{1}$ and $G_{2}$, which is denoted by $G_{1}\cup G_{2}$. For any integer $k\geq 1$, let $kG$ denote the disjoint union of $k$ copies of $G$. The join of $G_{1}$ and $G_{2}$ is denoted by $G_{1}\vee G_{2}$, which is obtained from $G_{1}\cup G_{2}$ by joining each vertex of $G_{1}$ with each vertex of $G_{2}$ by an edge.
The binding number was first introduced by Woodall [1] and Anderson [2]. The binding number of $G$, denoted by $bind(G)$, is defined by
$bind(G)=\min\left\{\frac{|N_{G}(S)|}{|S|}:\emptyset\neq S\subseteq V(G),N_{G}(S)\neq V(G)\right\}.$
A graph $G$ is said to be $r$-binding if $bind(G)\geq r$, where $r$ is a positive real number. We refer the reader to [3, 4, 5, 6, 7] for some properties on binding numbers.
Let $G$ be a graph with $V(G)=\{v_{1},v_{2},\ldots,v_{n}\}$. The adjacency matrix of $G$ is the real symmetric matrix $A(G)=(a_{ij})_{n\times n}$, where $a_{ij}=1$ if $v_{i}v_{j}\in E(G)$ and $a_{ij}=0$ if $v_{i}v_{j}\notin E(G)$. The largest eigenvalue of $A(G)$, denoted by $\rho(G)$, is called the adjacency spectral radius of $G$. For some results on spectral radius in graphs, we refer the reader to [8, 9, 10, 11, 12, 13, 14, 15, 16].
Let $H:V(G)\longrightarrow\{\{1\},\{0,2\}\}$ be a set-valued function. Hence, $H(v)$ equals $\{1\}$ or $\{0,2\}$ for any $v\in V(G)$. We let
$H^{-1}(1)=\{v:v\in V(G)\ \mbox{and}\ H(v)=1\}.$
An $H$-factor of $G$ is a spanning subgraph $F$ of $G$ such that $d_{F}(v)\in H(v)$ for each $v\in V(G)$. In terms of the Handshaking Lemma, we easily see that $H^{-1}(1)$ is even if $G$ contains an $H$-factor. Obviously, $G$ contains no $H$-factor if $H^{-1}(1)$ is odd.
Lu and Kano [17] obtained a characterization for the existence of an $H$-factor in a graph.
Theorem 1.1 (Lu and Kano [17]). A connected graph $G$ contains an $H$-factor for each $H:V(G)\longrightarrow\{\{1\},\{0,2\}\}$ with $H^{-1}(1)$ even if and only if
$\omega(G-S)\leq|S|+1$
for any $S\subseteq V(G)$.
Motivated by [17, 18] directly, we establish a sufficient condition with respect to the adjacency spectral radius to guarantee the existence of an $H$-factor in a 1-binding graph.
Theorem 1.2. Let $G$ be a connected 1-binding graph of order $n\geq 11$. If
$\rho(G)\geq\rho(K_{1}\vee(K_{n-4}\cup K_{2}\cup K_{1})),$
then $G$ has an $H$-factor for each $H:V(G)\longrightarrow\{\{1\},\{0,2\}\}$ with $H^{-1}(1)$ even, unless $G=K_{1}\vee(K_{n-4}\cup K_{2}\cup K_{1})$.
2 Some preliminaries
In this section, we provide several necessary preliminary lemmas, which will be used to verify our main result.
Lemma 2.1 (Li and Feng [19]). Let $G$ denote a graph, and let $H$ be a subgraph of $G$. Then
$\rho(G)\geq\rho(H),$
with equality occurring if and only if $G=H$.
Lemma 2.2. Let $\sum\limits_{i=1}^{t}n_{i}=n-s$ with $s\geq 1$. If $n_{t}\geq n_{t-1}\geq\cdots\geq n_{1}\geq 1$, $n_{t-1}\geq p\geq 1$ and $n_{t}<n-s-t-p+2$, then
$\rho(K_{s}\vee(K_{n_{t}}\cup K_{n_{t-1}}\cup\cdots\cup K_{n_{1}}))<\rho(K_{s}\vee(K_{n-s-t-p+2}\cup K_{p}\cup(t-2)K_{1})).$
Proof. Write $G=K_{s}\vee(K_{n_{t}}\cup K_{n_{t-1}}\cup\cdots\cup K_{n_{1}})$, and let $\mathbf{x}$ be the Perron vector of $A(G)$. Using symmetry, we let $\mathbf{x}(v)=x_{i}$ for any $v\in V(K_{n_{i}})$, where $1\leq i\leq t$, and $\mathbf{x}(u)=y$ for any $u\in V(K_{s})$. Notice that $K_{n_{t}+s}$ is a proper subgraph of $G$. Together with Lemma 2.1, we deduce $\rho(G)>\rho(K_{n_{t}+s})=n_{t}+s-1>n_{t}-1$. Notice that $n_{t}\geq n_{t-1}\geq\cdots\geq n_{1}\geq 1$. According to $A(G)\mathbf{x}=\rho(G)\mathbf{x}$, we possess
$(\rho(G)-(n_{i}-1))(x_{t}-x_{i})=$ $(\rho(G)-(n_{i}-1))x_{t}-(\rho(G)-(n_{i}-1))x_{i}$
$=$ $(\rho(G)-(n_{i}-1))x_{t}-sy$
$=$ $(\rho(G)-(n_{t}-1)+(n_{t}-1)-(n_{i}-1))x_{t}-sy$
$=$ $sy+(n_{t}-n_{i})x_{t}-sy$
$=$ $(n_{t}-n_{i})x_{t}$
$\geq$ 00 0
for $1\leq i\leq t$. This leads to $x_{t}\geq x_{i}$ for $1\leq i\leq t$. Write $G^{\prime}=K_{s}\vee(K_{n-s-t-p+2}\cup K_{p}\cup(t-2)K_{1})$. Then it follows from $p\leq n_{t-1}\leq n_{t}<n-s-t-p+2$, $n_{j}\geq 1$ for $21\leq j\leq t-$ 2 and $x_{t}\geq x_{i}$ for $1\leq i\leq t$ that
$\rho(G^{\prime})-\rho(G)\geq$ $\mathbf{x}^{T}(A(G^{\prime})-A(G))\mathbf{x}$
$=$ $2(n_{t-1}-p)x_{t-1}(n_{t}x_{t}+\sum\limits_{j=1}^{t-2}(n_{j}-1)x_{j}-px_{t-1})$
$+2\sum\limits_{j=1}^{t-2}(n_{j}-1)x_{j}(n_{t}x_{t}-x_{j})$
$+2\sum\limits_{i=2}^{t-2}\sum\limits_{j=i-1}^{t-3}(n_{i}-1)(n_{j}-1)x_{i}x_{j}$
$>$ $0,$
that is, $\rho(K_{s}\vee(K_{n_{t}}\cup K_{n_{t-1}}\cup\cdots\cup K_{n_{1}}))<\rho(K_{s}\vee(K_{n-s-t-p+2}\cup K_{p}\cup(t-2)K_{1}))$. This completes the proof of Lemma 2.2. $\Box$
Let $M$ denote a real matrix of order $n$ and $\mathcal{N}=\{1,2,\ldots,n\}$. For a given partition $\pi:\mathcal{N}=\mathcal{N}_{1}\cup\mathcal{N}_{2}\cup\cdots\cup\mathcal{N}_{r}$, the matrix $M$ can be correspondingly written as
$M=\left(\begin{array}[]{cccc}M_{11}&M_{12}&\cdots&M_{1r}\\ M_{21}&M_{22}&\cdots&M_{2r}\\ \vdots&\vdots&\ddots&\vdots\\ M_{r1}&M_{r2}&\cdots&M_{rr}\\ \end{array}\right),$
where the block $M_{ij}$ denotes the $n_{i}\times n_{j}$ matrix for any $1\leq i,j\leq r$. Let $m_{ij}$ denote the average row sum of $M_{ij}$, that is to say, $m_{ij}$ is the sum of all entries in $M_{ij}$ divided by the number of rows. Then the $r\times r$ matrix $M_{\pi}=(m_{ij})$ is called the quotient matrix of $M$. The partition $\pi$ is called equitable if every block $M_{ij}$ of $M$ admits constant row sum $m_{ij}$ for $1\leq i,j\leq r$.
Lemma 2.3 (Brouwer and Haemers [20], You, Yang, So and Xi [21]). Let $M$ be a real matrix of order $n$ with an equitable partition $\pi$, and let $M_{\pi}$ be the corresponding quotient matrix. Then the eigenvalues of $M_{\pi}$ are eigenvalues of $M$. Furthermore, if $M$ is nonnegative and irreducible, then the largest eigenvalues of $M$ and $M_{\pi}$ are equal.
3 Proof of Theorem 1.2
Proof of Theorem 1.2. Suppose that a connected 1-binding graph $G$ has no $H$-factor. Using Theorem 1.1, we conclude $\omega(G-S)\geq|S|+2$ for some nonempty subset $S$ of $V(G)$. Let $|S|=s$ and $\omega(G-S)=q$. Then $q\geq s+2$. Let $C_{1},C_{2},\ldots,C_{q-1},C_{q}$ denote the $q$ components in $G-S$ and $|C_{i}|=n_{i}$ for $1\leq i\leq q$. Without loss of generality, we let $n_{q}\geq n_{q-1}\geq\cdots\geq n_{2}\geq n_{1}\geq 1$. We are to verify $n_{s+1}\geq 2$. Otherwise, if $n_{i}=1$ for $11\leq i\leq s+$ 1, then we let $X=V(C_{1})\cup V(C_{2})\cup\cdots\cup V(C_{s+1})$. Obviously, $N_{G}(X)\subseteq S$ and $|X|=s+1$. Thus, we deduce
$\frac{|N_{G}(X)|}{|X|}\leq\frac{|S|}{|X|}=\frac{s}{s+1}<1,$
which is impossible because $G$ is 1-binding. This leads to $n_{i}\geq 2$ for $i\geq s+1$. We easily see that $G$ is a spanning subgraph of $G_{1}=K_{s}\vee(K_{n_{1}}\cup K_{n_{2}}\cup\cdots\cup K_{n_{s+1}}\cup K_{n_{s+2}})$, where $n_{s+2}\geq n_{s+1}\geq\cdots\geq n_{2}\geq n_{1}\geq 1$, $n_{s+1}\geq 2$ and $\sum\limits_{i=1}^{s+2}n_{i}=n-s$. In view of Lemma 2.1, we get
$\rho(G)\leq\rho(G_{1}),$ (3.1)
with equality occurring if and only if $G=G_{1}$. Let $G_{2}=K_{s}\vee(K_{n-2s-2}\cup K_{2}\cup sK_{1})$, where $n\geq 2s+4$. According to Lemma 2.2 (set $p=2$ and $t=s+2$ in Lemma 2.2), we conclude
$\rho(G_{1})\leq\rho(G_{2}),$ (3.2)
where the equality occurs if and only if $G_{1}=G_{2}$. For $s=1$, we admit $G_{2}=K_{1}\vee(K_{n-4}\cup K_{2}\cup K_{1})$. Using (3.1) and (3.2), we infer
$\rho(G)\leq\rho(K_{1}\vee(K_{n-4}\cup K_{2}\cup K_{1})),$
with equality if and only if $G=K_{1}\vee(K_{n-4}\cup K_{2}\cup K_{1})$. Observe that $K_{1}\vee(K_{n-4}\cup K_{2}\cup K_{1})$ contains no $H$-factor, a contradiction. We are to consider $s\geq 2$
Case 1. $n\geq 2s+5$.
Recall that $G_{2}=K_{s}\vee(K_{n-2s-2}\cup K_{2}\cup sK_{1})$. For the partition $V(G_{2})=V(K_{s})\cup V(K_{n-2s-2})\cup V(K_{2})\cup V(sK_{1})$, the quotient matrix of $A(G_{2})$ can be written as
$B_{2}=\left(\begin{array}[]{cccc}s-1&n-2s-2&2&s\\ s&n-2s-3&0&0\\ s&0&1&0\\ s&0&0&0\\ \end{array}\right).$
Hence, the characteristic polynomial of $B_{2}$ equals
$\varphi_{B_{2}}(x)=$ $x^{4}-(n-s-3)x^{3}-(s^{2}+2s+1)x^{2}+(s^{2}n+2sn+n-2s^{3}-6s^{2}-7s-3)x$
$-s^{2}n+2s^{3}+3s^{2}.$
One easily see that the partition $V(G_{2})=V(K_{s})\cup V(K_{n-2s-2})\cup V(K_{2})\cup V(sK_{1})$ is equitable. From Lemma 2.3, $\rho(G_{2})$ is the largest root of $\varphi_{B_{2}}(x)=0$.
Let $G_{*}=K_{1}\vee(K_{n-4}\cup K_{2}\cup K_{1})$. By the partition $V(G_{*})=V(K_{1})\cup V(K_{n-4})\cup V(K_{2})\cup V(K_{1})$, the quotient matrix of $A(G_{*})$ is given by
$B_{*}=\left(\begin{array}[]{cccc}0&n-4&2&1\\ 1&n-5&0&0\\ 1&0&1&0\\ 1&0&0&0\\ \end{array}\right),$
and so its characteristic polynomial is
$\varphi_{B_{*}}(x)=x^{4}-(n-4)x^{3}-4x^{2}+(4n-18)x-n+5.$
Since the partition $V(G_{*})=V(K_{1})\cup V(K_{n-4})\cup V(K_{2})\cup V(K_{1})$ is equitable, it follows from Lemma 2.3 that $\rho(G_{*})$ is the largest root of $\varphi_{B_{*}}(x)=0$.
Note that $\varphi_{B_{2}}(\rho(G_{2}))=0$. Thus, we possess
$\varphi_{B_{*}}(\rho(G_{2}))=\varphi_{B_{*}}(\rho(G_{2}))-\varphi_{B_{2}}(\rho(G_{2}))=(s-1)f(\rho(G_{2})),$ (3.3)
where $f(\rho(G_{2}))=-\rho^{3}(G_{2})+(s+3)\rho^{2}(G_{2})-((s+3)n-2s^{2}-8s-15)\rho(G_{2})+(s+1)n-2s^{2}-5s-5$.
Since $K_{n-s-2}$ is a proper subgraph of $G_{2}=K_{s}\vee(K_{n-2s-2}\cup K_{2}\cup sK_{1})$, it follows from Lemma 2.1 that
$\rho(G_{2})>\rho(K_{n-s-2})=n-s-3.$ (3.4)
Let $f(x)=-x^{3}+(s+3)x^{2}-((s+3)n-2s^{2}-8s-15)x+(s+1)n-2s^{2}-5s-5$, $x\geq n-s-3$. Then the derivative function of $f(x)$ is
$f^{\prime}(x)=-3x^{2}+2(s+3)x-(s+3)n+2s^{2}+8s+15,\ \ \ \ \ x\geq n-s-3.$
Notice that $\frac{s+3}{3}<n-s-3\leq x$ by $s\geq 2$ and $n\geq 2s+5$. Thus, we infer
$f^{\prime}(x)\leq$ $f^{\prime}(n-s-3)$
$=$ $-3n^{2}+7(s+3)n-3s^{2}-22s-30$
$\leq$ $-3(2s+5)^{2}+7(s+3)(2s+5)-3s^{2}-22s-30\ \ \ \ \ \ \ (\mbox{since}\ n\geq 2s+5)$
$=$ $-s^{2}-5s$
$<$ $0$
This implies that $f(x)$ is decreasing with respect to $x\geq n-s-3$. Combining this with (3.4), we deduce
$f(\rho(G_{2}))<$ $f(n-s-3)$
$=$ $-(n-s-3)^{3}+(s+3)(n-s-3)^{2}$
$-((s+3)n-2s^{2}-8s-15)(n-s-3)+(s+1)n-2s^{2}-5s-5$
$=$ $-n^{3}+(3s+9)n^{2}-(2s^{2}+15s+20)n+2s^{2}+10s+4.$ (3.5)
Let $g(n)=-n^{3}+(3s+9)n^{2}-(2s^{2}+15s+20)n+2s^{2}+10s+4$. Using a similar discussion as above, we easily see that $g(n)$ is decreasing with respect to $n\geq 2s+5$. In terms of $s\geq 2$ and $n\geq 2s+5$, we get
$g(n)\leq$ $g(2s+5)$
$=$ $-(2s+5)^{3}+(3s+9)(2s+5)^{2}-(2s^{2}+15s+20)(2s+5)+2s^{2}+10s+4$
$=$ $-2s^{2}+4$
$<$ $0$
Combining this with (3.3), (3) and $s\geq 2$, we obtain
$\varphi_{B_{*}}(\rho(G_{2}))=(s-1)f(\rho(G_{2}))<(s-1)g(n)<0,$
which yields that
$\rho(G_{2})<\rho(G_{*}).$ (3.6)
By virtue of (3.1), (3.2) and (3.6), we conclude
$\rho(G)\leq\rho(G_{1})\leq\rho(G_{2})<\rho(G_{*})=\rho(K_{1}\vee(K_{n-4}\cup K_{2}\cup K_{1})),$
which contradicts $\rho(G)\geq\rho(K_{1}\vee(K_{n-4}\cup K_{2}\cup K_{1}))$.
Case 2. $n=2s+4$.
In this case, $G_{2}=K_{s}\vee(2K_{2}\cup sK_{1})$. Consider the partition $V(G_{2})=V(K_{s})\cup V(2K_{2})\cup V(sK_{1})$, the corresponding quotient matrix of $A(G_{2})$ is
$B_{3}=\left(\begin{array}[]{ccc}s-1&4&s\\ s&1&0\\ s&0&0\\ \end{array}\right).$
The characteristic polynomial of $B_{3}$ equals
$\varphi_{B_{3}}(x)=x^{3}-sx^{2}-(s^{2}+3s+1)x+s^{2}.$
Note that the partition $V(G_{2})=V(K_{s})\cup V(2K_{2})\cup V(sK_{1})$ is equitable. By Lemma 2.3, $\rho(G_{2})$ is the largest root of $\varphi_{B_{3}}(x)=0$. That is, $\varphi_{B_{3}}(\rho(G_{2}))=0$.
If $s\in\{2,3\}$, then $n\in\{8,10\}$, which contradicts $n\geq 11$. Next, we discuss $s\geq 4$.
Since $K_{s}\vee(s+4)K_{1}$ is a proper subgraph of $G_{2}=K_{s}\vee(2K_{2}\cup sK_{1})$, it follows from Lemma 2.1 that
$\rho(G_{2})>\rho(K_{s}\vee(s+4)K_{1}).$ (3.7)
In terms of the partition $V(K_{s}\vee(s+4)K_{1})=V(K_{s})\cup V((s+4)K_{1})$, the quotient matrix of $A(K_{s}\vee(s+4)K_{1})$ equals
$B_{4}=\left(\begin{array}[]{ccc}s-1&s+4\\ s&0\\ \end{array}\right).$
The characteristic polynomial of $B_{4}$ is
$\varphi_{B_{4}}(x)=x^{2}-(s-1)x-s(s+4).$
Note that the partition $V(K_{s}\vee(s+4)K_{1})=V(K_{s})\cup V((s+4)K_{1})$ is equitable. According to Lemma 2.3, $\rho(K_{s}\vee(s+4)K_{1})$ is the largest root of $\varphi_{B_{4}}(x)=0$. Together with (3.7) and $s\geq 2$, we conclude
$\rho(G_{2})>\rho(K_{s}\vee(s+4)K_{1})=\frac{s-1+\sqrt{5s^{2}+14s+1}}{2}\geq\frac{3}{2}s+1.$ (3.8)
Recall that $n=2s+4$, $\varphi_{B_{*}}(x)=x^{4}-(n-4)x^{3}-4x^{2}+(4n-18)x-n+5$, $\varphi_{B_{3}}(x)=x^{3}-sx^{2}-(s^{2}+3s+1)x+s^{2}$ and $\varphi_{B_{3}}(\rho(G_{2}))=0$. Thus, we have
$\varphi_{B_{*}}(\rho(G_{2}))=$ $\varphi_{B_{*}}(\rho(G_{2}))-\rho(G_{2})\varphi_{B_{3}}(\rho(G_{2}))$
$=$ $-s\rho^{3}(G_{2})+(s^{2}+3s-3)\rho^{2}(G_{2})-(s^{2}-8s+2)\rho(G_{2})-2s+1.$ (3.9)
Let $h(x)=-sx^{3}+(s^{2}+3s-3)x^{2}-(s^{2}-8s+2)x-2s+1$, $x\geq\frac{3}{2}s+1$. Then the derivative function of $h(x)$ is
$h^{\prime}(x)=-3sx^{2}+2(s^{2}+3s-3)x-s^{2}+8s-2,\ \ \ \ \ x\geq\frac{3}{2}s+1.$
Since $s\geq 4$ and $\frac{2(s^{2}+3s-3)}{6s}<\frac{3}{2}s+1\leq x$, we possess
$h^{\prime}(x)\leq$ $h^{\prime}\Big{(}\frac{3}{2}s+1\Big{)}$
$=$ $-3s\Big{(}\frac{3}{2}s+1\Big{)}^{2}+2(s^{2}+3s-3)\Big{(}\frac{3}{2}s+1\Big{)}-s^{2}+8s-2$
$=$ $-\frac{15}{4}s^{3}+s^{2}-\frac{5}{2}s-11$
$<$ $0,$
which implies that $h(x)$ is decreasing with respect to $x\geq\frac{3}{2}s+1$. By (3.8) and $s\geq 4$, we obtain
$h(\rho(G_{2}))<$ $h\Big{(}\frac{s-1+\sqrt{5s^{2}+14s+1}}{2}\Big{)}$
$=$ $-s\Big{(}\frac{s-1+\sqrt{5s^{2}+14s+1}}{2}\Big{)}^{3}+(s^{2}+3s-3)\Big{(}\frac{s-1+\sqrt{5s^{2}+14s+1}}{2}\Big{)}^{2}$
$-(s^{2}-8s+2)\Big{(}\frac{s-1+\sqrt{5s^{2}+14s+1}}{2}\Big{)}-2s+1$
$=$ $\frac{1}{2}(-s^{4}+8s^{3}+28s^{2}-28s+3-(s^{3}+s^{2}-s-1)\sqrt{5s^{2}+14s+1})-\frac{1}{4}$
$<$ $\frac{1}{2}(-s^{4}+8s^{3}+28s^{2}-28s+3-(s^{3}+s^{2}-s-1)(2s+3))$
$=$ $\frac{1}{2}(-3s^{4}+3s^{3}+27s^{2}-23s+6)$
$<$ $0$
Combining this with (3), we deduce
$\varphi_{B_{*}}(\rho(G_{2}))=h(\rho(G_{2}))<0,$
which leads to $\rho(G_{2})<\rho(G_{*})$. Together with (3.1) and (3.2), we conclude
$\rho(G)\leq\rho(G_{1})\leq\rho(G_{2})<\rho(G_{*})=\rho(K_{1}\vee(K_{n-4}\cup K_{2}\cup K_{1})),$
which contradicts $\rho(G)\geq\rho(K_{1}\vee(K_{n-4}\cup K_{2}\cup K_{1}))$. Theorem 1.2 is proved. $\Box$
Data availability statement
My manuscript has no associated data.
Declaration of competing interest
The authors declare that they have no conflicts of interest to this work.
Acknowledgments
This work was supported by the Natural Science Foundation of Jiangsu Province (Grant No. BK20241949). Project ZR2023MA078 supported by Shandong Provincial Natural Science Foundation.

### v2

@@
 For vertex-disjoint graphs $G_{1}$ and $G_{2}$, the graph with vertex set $V(G_{1})\cup V(G_{2})$ and edge set $E(G_{1})\cup E(G_{2})$ is called the union of $G_{1}$ and $G_{2}$, which is denoted by $G_{1}\cup G_{2}$.
-For any integer $k\geq 1$, let $kG$ denote the disjoint union of $k$ copies of $G$.
+For any integer $k\geq 1$, let kGkG denote the disjoint union of $k$ copies of $G$.
 The join of $G_{1}$ and $G_{2}$ is denoted by $G_{1}\vee G_{2}$, which is obtained from $G_{1}\cup G_{2}$ by joining each vertex of $G_{1}$ with each vertex of $G_{2}$ by an edge.
@@
 $bind(G)=\min\left\{\frac{|N_{G}(S)|}{|S|}:\emptyset\neq S\subseteq V(G),N_{G}(S)\neq V(G)\right\}.$
-A graph $G$ is said to be $r$-binding if $bind(G)\geq r$, where $r$ is a positive real number. We refer the reader to [3, 4, 5, 6, 7] for some properties on binding numbers.
+A graph $G$ is said to be $r$-binding if $bind(G)\geq r$, where $r$ is a positive real number. We refer the reader to [3, 4, 5, 6, 7, 8] for some properties on binding numbers.
 Let $G$ be a graph with $V(G)=\{v_{1},v_{2},\ldots,v_{n}\}$.
@@
 The largest eigenvalue of $A(G)$, denoted by $\rho(G)$, is called the adjacency spectral radius of $G$.
-For some results on spectral radius in graphs, we refer the reader to [8, 9, 10, 11, 12, 13, 14, 15, 16].
+For some results on spectral radius in graphs, we refer the reader to [9, 10, 11, 12, 13, 14, 15, 16, 17, 18].
 Let $H:V(G)\longrightarrow\{\{1\},\{0,2\}\}$ be a set-valued function. Hence, $H(v)$ equals $\{1\}$ or $\{0,2\}$ for any $v\in V(G)$. We let
@@
 with equality occurring if and only if $G=H$.
-Lemma 2.2. Let $\sum\limits_{i=1}^{t}n_{i}=n-s$ with $s\geq 1$. If $n_{t}\geq n_{t-1}\geq\cdots\geq n_{1}\geq 1$, $n_{t-1}\geq p\geq 1$ and $n_{t}<n-s-t-p+2$, then
+Lemma $2.$ Let $\sum\limits_{i=1}^{t}n_{i}=n-s$ with $s\geq 1$. If $n_{t}\geq n_{t-1}\geq\cdots\geq n_{1}\geq 1$, $n_{t-1}\geq p\geq 1$ and $n_{t}<n-s-t-p+2$, then
 $\rho(K_{s}\vee(K_{n_{t}}\cup K_{n_{t-1}}\cup\cdots\cup K_{n_{1}}))<\rho(K_{s}\vee(K_{n-s-t-p+2}\cup K_{p}\cup(t-2)K_{1})).$
@@
 $=$ $(n_{t}-n_{i})x_{t}$
-$\geq$ 00 0
+$\geq$ $0$
 for $1\leq i\leq t$.
@@
 Write $G^{\prime}=K_{s}\vee(K_{n-s-t-p+2}\cup K_{p}\cup(t-2)K_{1})$.
-Then it follows from $p\leq n_{t-1}\leq n_{t}<n-s-t-p+2$, $n_{j}\geq 1$ for $21\leq j\leq t-$ 2 and $x_{t}\geq x_{i}$ for $1\leq i\leq t$ that
+Then it follows from $p\leq n_{t-1}\leq n_{t}<n-s-t-p+2$, $n_{j}\geq 1$ for $1\leq j\leq t-2$ and $x_{t}\geq x_{i}$ for $1\leq i\leq t$ that
 $\rho(G^{\prime})-\rho(G)\geq$ $\mathbf{x}^{T}(A(G^{\prime})-A(G))\mathbf{x}$
@@
 $>$ $0,$
-that is, $\rho(K_{s}\vee(K_{n_{t}}\cup K_{n_{t-1}}\cup\cdots\cup K_{n_{1}}))<\rho(K_{s}\vee(K_{n-s-t-p+2}\cup K_{p}\cup(t-2)K_{1}))$. This completes the proof of Lemma 2.2. $\Box$
+that is, $\rho(K_{s}\vee(K_{n_{t}}\cup K_{n_{t-1}}\cup\cdots\cup K_{n_{1}}))<\rho(K_{s}\vee(K_{n-s-t-p+2}\cup K_{p}\cup(t-2)K_{1}))$. This completes the proof of Lemma $2.$ $\Box$
 Let $M$ denote a real matrix of order $n$ and $\mathcal{N}=\{1,2,\ldots,n\}$.
@@
 Then $q\geq s+2$.
-Let $C_{1},C_{2},\ldots,C_{q-1},C_{q}$ denote the $q$ components in $G-S$ and $|C_{i}|=n_{i}$ for $1\leq i\leq q$.
+Let $C_{1},C_{2},\ldots,C_{q-1},C_{q}$ denote the $q$ components in GSG-S and $|C_{i}|=n_{i}$ for $1\leq i\leq q$.
 Without loss of generality, we let $n_{q}\geq n_{q-1}\geq\cdots\geq n_{2}\geq n_{1}\geq 1$.
 We are to verify $n_{s+1}\geq 2$.
-Otherwise, if $n_{i}=1$ for $11\leq i\leq s+$ 1, then we let $X=V(C_{1})\cup V(C_{2})\cup\cdots\cup V(C_{s+1})$.
+Otherwise, if $n_{i}=1$ for $1\leq i\leq s+1$, then we let $X=V(C_{1})\cup V(C_{2})\cup\cdots\cup V(C_{s+1})$.
 Obviously, $N_{G}(X)\subseteq S$ and $|X|=s+1$.
@@
 $=$ $-s^{2}-5s$
-$<$ $0$
+$<$ $0.$
 This implies that $f(x)$ is decreasing with respect to $x\geq n-s-3$. Combining this with (3.4), we deduce
@@
 $=$ $-2s^{2}+4$
-$<$ $0$
+$<$ $0.$
 Combining this with (3.3), (3) and $s\geq 2$, we obtain
@@
 $=$ $\frac{1}{2}(-3s^{4}+3s^{3}+27s^{2}-23s+6)$
-$<$ $0$
+$<$ $0.$
 Combining this with (3), we deduce